Amina thinks of a number and subtracts 5/2 from it. She multiplies the result by 8. the final result is 3 times her original number. Find the number.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the number thought by Amina be x

She subtracts 5/2 from the number.

x – 5/2

Now, she multiplies the result by 8

8(x – 5/2)

The result obtained is three times her original number;

8(x – 5/2) = 3x

8(x – 5/2) = 3x

8x – 20 = 3x

5x = 20

x = 4

Hence, the number is 4.

Thanks